OpenStudy (anonymous):

Hint: rationalize the denominator. lim(x,y)→(0,0) for (−9x2+8y2) / ((−9x2+8y2+1)-1)

OpenStudy (anonymous):

so many parentheses D: If you know calculus you could just utilize L'Hopital's rule. Otherwise I suppose you'll just have to follow the hint.

OpenStudy (turingtest):

\[\lim_{(x,y) \rightarrow (0,0)}{-9x^2+8y^2 \over (-9x^2+8y^2+1)-1}\]right?

OpenStudy (anonymous):

I have to find the limit, and I am not sure how to rationalize the denom

OpenStudy (anonymous):

You have to multiply by the conjugate.

OpenStudy (anonymous):

To:TuringTest Thats correct

OpenStudy (turingtest):

why can't we just get rid of that 1 in the denominator and say\[\lim_{(x,y) \rightarrow (0,0)}{-9x^2+8y^2\over-9x^2+8y^2}=1\]

OpenStudy (turingtest):

I don't see how to rationalize that denominator. It doesn't look irrational to me.

OpenStudy (anonymous):

Thats my problem It think it may want polar cordinates, and idk how to do those either

OpenStudy (anonymous):

if we do your method, wouldnt it just turn into 1/(1-1) making it irrational

OpenStudy (turingtest):

How did you get a 1 in the numerator? as far as the denominator, look at what I did above. You can ignore those parentheses, they are meaningless. So if you rewrite the denominator as\[-9x^2+8y^2+1-1=-9x^2+8y^2\]you see it's the same as the numerator. You clearly don't need to plug in (0,0) when the top and bottom cancel to give 1.

OpenStudy (anonymous):

Thats wrong I submitted 1 and it came up incorrect Im an idiot, I type the problem wrong! The part in parenthese is sqrt's that is why the neg 1 is on the outside

OpenStudy (turingtest):

Ahhh, that explains everything. It also makes the problem much more difficult. Let me see now...\[\lim_{(x,y) \rightarrow (0,0)}={-9x^2+8x^2\over \sqrt{-9x^2+8y^2+1}-1}\]right? Hmm... let me try some things on paper, it'll be faster that way. I'll have to get back to you (hopefully not too long).

OpenStudy (anonymous):

Im so sorry for the confusion!

OpenStudy (turingtest):

it's all good... working on it...

OpenStudy (turingtest):

ok it's not so bad, check it out... after multiplying top and bottom by the conjugate we get\[{(-9x^2+8x^2)(\sqrt{-9x^2+8y^2+1}+1)\over (-9x^2+8x^2)}=\sqrt{-9x^2+8y^2+1}+1\]so\[\lim_{(x,y) \rightarrow (0,0)}\sqrt{-9x^2+8y^2+1}+1=2\]I hope you see how I got the denominator after multiplying by the conjugate.

OpenStudy (anonymous):

Thats correct, wow good job

OpenStudy (anonymous):

I hate conjugates, but yeah I can see what you did

OpenStudy (turingtest):

Thanks, it was kinda tough. I don't usually beg but... medal please? I took a bit of time on this one and I think I earned it.

OpenStudy (turingtest):

thx

OpenStudy (anonymous):

dude yeah!! I give medals to all who help and show effort I can most more similar problems if you want more medals

OpenStudy (anonymous):

none will be as hard im sure, and i dont need work

OpenStudy (anonymous):

i just need the answer haha

OpenStudy (turingtest):

I'm getting drowsy, going to sleep soon. However most tutors here can't do these so if I can't do it tonight I'll look up your questions tomorrow. If you think they're easier though and you just need answers you can try me. Just post separately.

OpenStudy (anonymous):

How do i post separately, can I direct them to you some how?

OpenStudy (turingtest):

Not that I know of, but I watch for the most recent questions. Just post it normally and I should find it.

OpenStudy (anonymous):

ok thanks!

OpenStudy (turingtest):

oh, and if you become my fan I can find your questions more easily when you are online, so you may want to do that.